### v1

@@ Abstract
Enterprises are currently undergoing profound transformations due to the unpostponable digital transformation. Then, to remain competitive, enterprises must adapt their organisational structures and operations. This organisational shift is also important for small and medium-sized enterprises. A key innovation frontier is the adoption of process-oriented production models. This paper presents a knowledge-based method to support business experts in designing business processes. The method requires no prior expertise in Knowledge Engineering and guides designers through a structured sequence of steps to produce a diagrammatic workflow of the target process. The construction of the knowledge base starts from simple, text-based, knowledge artefacts and then progresses towards more structured, formal representations. The approach has been conceived to allow a shared approach for all stakeholders and actors who participate in the BP design.
keywords:
Small-Medium Enterprises, Business Process Analysis Canvas, Knowledge-based Method, Cooperative Work Environment
@@ 2 Related Work
A variety of approaches have been proposed to support Business Process Analysis (BPA) through digital innovations and automation. Goni et al. examined sustainable business models and proposed a framework for embedding sustainability within business processes [4]. Ahmad and Van Looy identified key trends in the digital transformation of BPM, including strategic alignment and business-IT integration [5]. Schüler and Alpers categorised methods for automated process model generation and outlined challenges related to data quality and model applicability [6]. Similarly, Weinzierl et al. reviewed machine learning applications across the BPM lifecycle, highlighting their potential in optimisation and decision support [7].
In the area of Business Process Knowledge Bases (BPKBs), several frameworks have been developed to formalise and structure process knowledge. Motta et al. proposed a web-based knowledge management system that combines performance metrics, process structuring, and benchmarking [8]. Missikoff introduced a methodology for BPA based on Model-Driven Architecture, transforming textual process descriptions into ontologies [9]. Ligeza and Potempa combined BPMN with business rules and logic-based models to enhance semantic accuracy and automate process execution [10].
@@ 3 The BPA Canvas Methodology
In this section, we present the BPA Canvas methodology, that in its full representation is organised in eight knowledge segments. The first objective of the BPA Canvas is to facilitate the collection of the knowledge about a given BP in a collaborative, shared fashion. In this way, we ensure a large involvement of all the stakeholders and players that collaboratively participate in the specification of the BP. Such a wide participation contributes to improving the quality of the analysis report, reducing missing information, ambiguities and imprecision. The BPA Canvas also represents the structure of the shared BP knowledge base (BPKB), a repository structured accordingly. A collection of BPKBs are able to model a decentralised production ecosystem and the set of interconnected BPs taking place therein.
@@
    BP Signature. The first knowledge artefact, in the form of a table, aimed at providing a synthetic description of the business process.
@@
The collected knowledge is finally used to draw the BP workflow, i.e., a diagram representation that includes actors, tasks, and their sequencing. To this end we adopt the BPMN graphical standard [11].
@@ 4 Application of the BPA Canvas in MERA
MERA is the institution responsible for issuing the necessary licences to enterprises that want a connection to the electricity grid in Malawi. The licensing process may concern also a group of SMEs that wants to coordinate to obtain a collective licence. The primary objective of this process is to ensure that licences are granted solely to qualified and compliant individuals or enterprises. This approach safeguards the integrity, compliance, and safety standards within Malawi’s energy sector. The meticulous and systematic design of the licensing process ensures all regulatory requirements are fulfilled prior to licence issuance. In essence, the process starts when an applicant submits to MERA the form requesting the access to the electricity grid. The application form is verified to check the validity of the provided data and then an additional level of verification consists in the interview of the applicant. Finally, when all the data are complete and correct, the applicant is invoiced and, upon payment, the licence is issued.
@@ 4.1 BP Signature
The BP Signature represents the initial knowledge artefact in the BPA Canvas methodology. It provides a high-level overview of the business process by capturing its essential components in a concise, tabular format. Table 1 provides the BP signature corresponding to the MERA licensing process, detailing the principal actors, involved objects, input elements, outputs, and the main objective of the workflow.
@@
BP Name MERA Licence Application Process
Key Actors Applicant, MERA Team, Applicant, Licensing Office, Applicant, CEO
Key Objects Application Form, Application Form_Integration Request, Application Form_Integrated, Invoice, Licence Payment Receipt, Licence_Signed, Licence Registration Number, Application Submission Check List, Site Inspection Check List
@@ 4.2 BP Statement
    • The licensing process commences when an applicant submits a fully completed application form along with the required supporting documents to the Licensing Office.
    • Upon receipt, the Licensing Office registers the application and forwards it to the MERA Team for further review.
@@
    • The Board Committee conducts an interview with the applicant to evaluate their qualifications and capabilities relevant to the requested licence.
@@
    • Following payment verification, the application is forwarded to the CEO for final approval and signature.
@@ 4.3 BP User story
    After ensuring that the application is complete, Mr. Banda submits the form and documents to the Licensing Office. The Licensing Office receives his application, logs it into their system, and assigns it a reference number. The application is then forwarded to the MERA Team for further review.
@@
    The team promptly informs Mr. Banda of the missing information and requests that he provides the missing competencies. Mr. Banda quickly gathers the required data and submits it to MERA, which then updates the application. With all data now complete, the application moves forward to the next stage.
@@ 5 The BP Diagram
The BP diagram is represented by using the Business Process Model and Notation (BPMN) graphical standard [11]. As illustrated in Figure 1, the BP is represented as a pool in which actors are represented as lanes, and tasks are connected by directed arcs to indicate the flow of activities. Different nodes can represent tasks (boxes with rounded corners), decision points (diamonds), or events (circles). For a more exhaustive introduction, please refer to the wide literature and tutorials.
Figure 1: BPMN representation of the MERA licence application process
In the diagram we see five lanes, corresponding to the key actors of the BP, namely: Applicant, Licensing Office, MERA Team, Board Committee, and CEO. The BP starts with the Applicant filling out of the request form and ends with the rejection, in case of incompetent Applicant, or a positive completion.
The BPMN diagram visually represents the key stages of the MERA licensing process as described in the previous BPA Canvas segments. It serves both as a formal representation and a practical tool for communicating the structure and the logic of the workflow to various stakeholders. By aligning the diagram with the BP Signature, Statement, and User Story, the model ensures coherence between textual and visual artefacts, thereby enhancing understanding, and the potential for future automation. This graphical abstraction also provides a foundation for the representation of multiple processes belonging to a number of partner SMEs that are part of a broader decentralised production ecosystem.
@@ 6 Conclusion
This study introduced a knowledge-based methodology, namely, the BPA Canvas, as a structured and cooperative framework for business process analysis and representation. By enabling a common representation of business processes and involving key stakeholders in the knowledge modelling phase, the methodology contributes to produce high quality process design, reduced ambiguity, and more informed design decision-making. Through the case study of the MERA licensing process, we demonstrated how the BPA Canvas can be practically applied to capture operational knowledge, align goals, representing the foundation for cooperative process design.

### v2

@@ Abstract
Enterprises are currently undergoing profound transformations due to the unpostponable digital transformation. Then, to remain competitive, enterprises must adapt digital solutions, transforming their organisational structures and operations. This organisational shift is also important for small and medium-sized enterprises. A key innovation frontier is the adoption of process-oriented production models. This paper presents a knowledge-based method to support business experts in designing business processes. The method requires no prior expertise in Knowledge Engineering and guides designers through a structured sequence of steps to produce a diagrammatic workflow of the target process. The construction of the knowledge base starts from simple, text-based, knowledge artefacts and then progresses towards more structured, formal representations. The approach has been conceived to allow a shared approach for all stakeholders and actors who participate in the BP design.
@@ 2 Related Work
A variety of approaches have been proposed to support Business Process Analysis (BPA) aiming at digital innovations and process automation. Goni et al. examined sustainable business models and proposed a framework for embedding sustainability within business processes [6]. Ahmad and Van Looy identified key trends in the digital transformation of Business Process Management (BPM), including strategic alignment and business-IT integration [1]. Schüler and Alpers categorised methods for automated process model generation and outlined challenges related to data quality and model applicability [10]. Similarly, Weinzierl et al. reviewed machine learning applications across the BPM lifecycle, highlighting their potential in optimisation and decision support [11].
In the area of Business Process Knowledge Bases (BPKBs), several frameworks have been developed to formalise and structure process knowledge. Motta et al. proposed a web-based knowledge management system that combines performance metrics, process structuring, and benchmarking [9]. Missikoff introduced a methodology for BPA based on Model-Driven Architecture, transforming textual process descriptions into ontologies [8]. Ligeza and Potempa introduced an Artificial Intelligence (AI) framework that combines Business Process Model and Notation (BPMN) [4] with Business Rules to enhance semantic accuracy [7]. Azarijafari et al. introduced an Agentic AI approach to BP Development, shifting from traditional task-oriented workflows to a goal-driven model [3].
@@ 3 The BPA Canvas Methodology
In this section, we present the BPA Canvas methodology that in its full representation is organised in eight knowledge segments. The first objective of the BPA Canvas is to facilitate the collection of the knowledge about a given BP in a collaborative, shared fashion. In this way, we ensure a large involvement of all the stakeholders and players that collaboratively participate in the specification of the BP. Such a wide participation contributes to improving the quality of the analysis report, reducing missing information, ambiguities and imprecision. The BPA Canvas also represents the structure of the shared BP knowledge base (BPKB), a repository organised accordingly. A collection of BPKBs are able to model a decentralised production ecosystem and the set of interconnected BPs taking place therein.
@@
    BP Signature. The first knowledge artefact, in the form of a simple table, aimed at providing a synthetic description of the business process.
@@
The collected knowledge is finally used to draw the BP workflow, i.e., a diagram representation that includes actors, tasks, and their sequencing. To this end we adopt the BPMN graphical standard.
@@ 4 Application of the BPA Canvas in MERA
MERA is the institution responsible for issuing the necessary licences to enterprises that want a connection to the electricity grid in Malawi. The licensing process may concern also a group of SMEs that wants to coordinate to obtain a collective licence. The primary objective of this process is to ensure that licences are granted solely to qualified and compliant individuals or enterprises. This approach safeguards the integrity, compliance, and safety standards within Malawi’s energy sector. The meticulous and systematic design of the licensing process ensures that all regulatory requirements are fulfilled prior to licence issuance. In essence, the process starts when an applicant submits to MERA the form requesting the access to the electricity grid. The application form is verified to check the validity of the provided data and then an additional level of verification consists in the interview of the applicant. Finally, when all the data are complete and correct, the applicant is invoiced and, upon payment, the licence is issued.
@@ 4.1 BP Signature
The BP Signature represents the initial knowledge artefact in the BPA Canvas methodology. It provides a high-level overview of the business process by capturing its essential components in a concise, tabular format. Table 1 provides the BP signature corresponding to the MERA licensing process, detailing the principal actors, involved objects, input elements, outputs, and the main objective of the business process.
@@
BP Name MERA Licence Application
Key Actors Applicant, MERA Team, Licensing Office, CEO
Key Objects Application Form, Application Form Integration Request, Application Form Integrated, Invoice, Licence Payment Receipt, Licence Signed, Licence Registration Number, Application Submission Check List, Site Inspection Check List
@@ 4.2 BP Statement
    • The licensing process commences when an applicant submits a fully filled-out application form along with the required supporting documents to the Licensing Office.
    • Upon receipt, the Licensing Office registers the application and forwards it to the MERA Team for administrative check.
@@
    • The Board Committee conducts an interview with the applicant to evaluate its qualifications and capabilities relevant to the requested licence.
@@
    • Following the payment verification, the application is forwarded to the CEO for final approval and signature.
@@ 4.3 BP User story
    After ensuring that the application is complete, Mr. Banda submits the form and documents to the Licensing Office. The Licensing Office receives his application, logs it into its system, and assigns it a reference number. The application is then forwarded to the MERA Team for further review.
@@
    The team promptly informs Mr. Banda of the missing information and requests that he provides the missing competencies. Mr. Banda quickly gathers the required data and submits them to MERA, which then updates the application. Being all data now complete, the application moves forward to the next stage.
@@ 5 The BP Diagram
The BP diagram is a graphical representation of the process described in the BP Statement. As illustrated in Figure 1, the BP is represented using a BPMN diagram as a pool in which actors are represented as lanes, and tasks are connected by directed arcs to indicate the flow of activities. Different nodes can represent tasks (boxes with rounded corners), decision points (diamonds), or events (circles). For a more exhaustive introduction, please refer to the wide literature and tutorials.
Figure 1: BPMN representation of the MERA licencing process
In the diagram we see five lanes, corresponding to the key actors of the BP, namely: Applicant, Licensing Office, MERA Team, Board Committee, and CEO. The BP starts with the Applicant filling out the request form and ends with the rejection, in case of incompetent Applicant, or a positive completion.
The BPMN diagram visually represents the key stages of the MERA licensing process as described in the previous BPA Canvas segments. It serves both as a rigorous representation and a practical tool for communicating the structure and the logic of the workflow to various stakeholders. By aligning the diagram with the BP Signature, Statement, and User Story, the model ensures coherence between textual and visual artefacts, thereby enhancing understanding, and the potential for future automation. This graphical abstraction also provides a foundation for the representation of multiple processes belonging to a number of partner SMEs that are part of a broader decentralised production ecosystem.
@@ 6 Conclusion
This study introduced a knowledge-based methodology, namely, the BPA Canvas, as a structured and cooperative framework for business process analysis and representation. By enabling a common representation of business processes and involving key stakeholders in the knowledge modelling phase, the methodology contributes to produce high quality process design, reduced ambiguity, and more informed decision-making. Through the case study of the MERA licensing process, we demonstrated how the BPA Canvas, which represents the foundation for cooperative process design, can be practically applied to capture operational knowledge and align goals.